# Busting_the_princelings_replication.bib Are Princelings Truly Busted? Evaluating Transaction Discounts in China’s Land Market

### v1

@@ Abstract
This paper narrowly replicates Chen and Kung’s 2019 paper ( 134(1): 185-226). Inspecting the data reveals that nearly one-third of the transactions (388,903 out of 1,208,621) are perfect duplicates of other rows, excluding the transaction number. Replicating the analysis on the data sans-duplicates yields a slightly smaller but still statistically significant princeling effect, robust across the regression results. Further analysis also reveals that coefficients interpreted as the effect of logarithm of area actually reflect the effect of scaled values of area; this paper also reinterprets and contextualizes these results in light of the true scaled values.
@@ 1 Introduction
The recent implosion of several leading Chinese real estate developers, such as Evergrande and Country Garden, has kindled interest in the mechanisms of the Chinese real estate market. Unique not only for its highly leveraged financing practices, the market is also a nexus of interactions between state-owned and private enterprises and has historically been plagued by corruption and cronyism, particularly on the land sale side. One often-cited paper on the topic is Chen and Kung (2019), titled “Busting the ‘Princelings’: The Campaign against Corruption in China’s Primary Land Market,” which explores the unique mechanism of Chinese cronyism in the land sale market. Specifically, they examine how “princeling” firms—those firms tied to family members of China’s top governing body, the Politburo—receive substantial price discounts on land parcels after controlling for a variety of location-based and transaction-level variables ([chen_busting_2019]). They also have significant findings relating princeling power, promotion likelihood, and magnitude of discount: more powerful princelings obtain larger discounts, and provincial party secretaries providing discounts were more likely to be promoted, with the likelihood increasing by the magnitude of the discount and the area of discounted land sold.
@@
Between containing detailed information about officials’ promotions and establishing firms’ state-ownership and princeling statuses, the paper’s replication data is also of popular interest and has over 11,600 downloads from its archive on the Harvard Dataverse.111The paper’s replication data can be found at https://doi.org/10.7910/DVN/XW6OJT.
@@
This paper replicates the results of Chen and Kung (2019, hereafter “CK”), accounting for cases of apparent duplicates and verifying data consistency. An initial inspection of CK’s data reveals that nearly one-third of the transactions (388,903 out of 1,208,621) are perfect duplicates of other rows, excluding the transaction number: these duplicates suggest, for instance, that the same buyer purchased two (or more) identical parcels of land with the same quality rating, area, and usage in the same city, month, and year, at the same price. Returning to the initial database and examining records of land sale transactions suggests that some of these duplicates correctly refer to identical yet distinct parcels sold at the same time, but many other duplicates erroneously reference the same parcel of land multiple times. Importantly, over one-half of the princeling parcels in the full dataset (9,815 out of 19,812) are duplicates. Replicating the analysis on the data, sans-duplicates, yields similar results to CK’s original findings: while the magnitude of discount for princelings is slightly smaller than that found using the original dataset, the princeling effect is still large and statistically significant across the regression results.
@@
Yet, further inspection reveals that when CK aggregate parcel areas to group parcels (for instance, summing data from transactions to calculate the total area of land sold to each firm in each year), these area measures are misinterpreted: while CK interpret the variable as the logarithm of area, it is actually area divided by 1,000,000. I thus reinterpret CK’s originally calculated coefficients as scaled area values rather than logged area values, with and without duplicates. While the princeling effect remains stable, its magnitude becomes quite large, suggesting that princelings on average purchase tens of thousands of additional square meters of land per year compared to their nonprinceling peers—areas double, triple, or even quadruple the median parcel size, depending on the specification. In the Appendix, I include calculations utilizing the corrected logarithm of area purchased; these results are strange and somewhat unstable, with the princeling effect not only changing sign and direction between tables, but also dropping sharply prior to the anticorruption campaign.
@@ 2 Apparent Duplicates in the Data
In their paper, CK construct a set of “princeling” firms from several online sources and match them to land transaction data obtained from the Ministry of Land and Resources. I focus on this latter data, extracted from the Land Transaction Monitoring System (www.landchina.com). As each municipality’s “bureau of land and resources is required to report each land transaction in their jurisdiction electronically on this website,” per the Law of Land Management (p. 199), the dataset captures all land transactions.222China eliminated the secondary market for land use rights (LURs) in 2004, meaning the State is the sole seller of land, doing so at the municipal level ([li_urban_2019]; [gyourko_land_2022]). The Land Transaction Monitoring System provides details on land transactions with increasing specificity in recent years, detailing everything from transaction details like total payment amount, sale method, buyer name, and total area of land sold, to building restrictions, like floor area ratio and greening ratio, to location features of the parcel, including its province, city, and district, as well as a qualitative description of its location.333CK note that the system also provides the address and postal code of the parcel, “a two-digit code of land usage… [and] a three-digit industry code of the buyer’s firm” (p. 200). These fields are no longer included on the Land Transaction Monitoring System website nor are they included in CK’s replication data, which makes determining the extent of duplicates more complicated; it is possible that these information fields were removed from the database sometime after CK collected the data. Alternatively, CK may have obtained a slightly more comprehensive version of the website’s data directly from the ministry.
@@
Examining CK’s dataset reveals several cases of apparent duplicates: specifically, after removing the transaction identifier, which was likely added by CK in data processing, 388,903 land transactions appear to be perfect duplicates. This suggests, in effect, that in the same month and year, a single buyer purchased identical parcels for the same price and use in the same city. Table LABEL:table_of_duplicates shows the extent and distribution of these duplicates. Of the 1,208,621 land transactions in CK’s data, 819,718 are unique, and 299,167 of these have at least one duplicate. While most transactions have only one duplicate, others have several—as many as 158.
Importantly, a considerable number of princeling firms are represented among the duplicates: of the 19,812 transactions that were conducted by princelings in the original data, over half of these are duplicates, with a significant share having three or more seemingly identical transactions.
| Observations | |||
| Frequency of occurrence | All | State-owned | Princeling |
| 1 | 520,551 | 181,571 | 4,414 |
| 2 | 218,001 | 75,640 | 3,479 |
| 3 | 79,366 | 27,473 | 1,554 |
| 4 | 581 | 193 | 81 |
| 5-10 | 966 | 319 | 201 |
| 10-20 | 167 | 59 | 48 |
| 20-30 | 48 | 10 | 21 |
| 40-158 | 38 | 15 | 17 |
| Number of unique rows | 819,718 | 285,280 | 9,815 |
| Number of total rows in Chen and Kung | 1,208,621 | 420,004 | 19,812 |
| Number of duplicate rows | 388,903 | 134,724 | 9,997 |
| Number of unique rows with duplicates | 299,167 | 103,709 | 5,401 |
In light of these results, there are a couple of possibilities: first, that there are truly identical neighboring parcels that are sold at the same time and price to a single buyer. In effect, the Ministry of Land could be subdividing a larger parcel and selling it as two (or more) smaller parcels. Specific usages could also dictate purchasing multiple nearby parcels of the same size and usage at the same time—for instance, land for building mobile phone towers.444Investigating the Ministry of Land’s database illustrates that this is in fact the case: telecommunications companies like China Mobile have purchased several parcels of the same area at the same price in the same city—in this case, the parcels were quite small (48 square meters), likely for telecommunications equipment. Second, it is possible that the duplicate rows are erroneous and that the database itself has multiple land transactions that refer to the same underlying land parcel.
@@
Examining CK’s replication data is not enough to confirm whether these duplicates refer to the same underlying land parcel, given they do not include each parcel’s specific address. Yet, the transaction details can be matched with the original data from the Ministry of Land’s database, and the additional information available on the website can be used to determine if multiple transaction listings refer to the same parcel. Specifically, after being identified in the database, each land transaction can be linked with a parcel number via the transaction’s original sale announcement (the “Transfer Announcement”), and one can confirm how many parcels of each size are up for sale in the area at the time. With this information, duplicate land transaction listings that refer to the same underlying parcel can be confirmed as erroneous duplicates.
@@
Looking at the detailed transaction records indicates that both hypotheses are true: while there are some cases where identical parcels were sold to a single buyer with the same features, many other cases appear to refer to one underlying parcel and are incorrect duplicates. Some of these are quite easy to recognize as duplicate listings of the same land transaction when referencing the Ministry of Land’s database. For instance, inspecting the database shows multiple cases where two separate transaction listings have the same electronic lookup number—a strong indicator of an incorrect duplicate.555The electronic lookup number is the database’s main identifier/search key for unique land transactions. Having two “separate” transactions with the same electronic lookup number is thus impossible; it can only be one transaction listed twice in error. Examining the Transfer Announcement for these parcels further confirms that only one parcel of the given size is listed for sale. Thus, there should only be one listing for the parcel in the Ministry of Land’s database—and in CK’s data.
@@
Other cases are slightly less clear but are still strongly indicative of incorrect duplicates. For instance, two transactions have different electronic lookup numbers, but investigating the Transfer Announcement suggests that both listings refer to the same underlying parcel.666 In many cases, examining the full transaction details of the two listings reveals minor discrepancies between them—not contradictions, but rather missing information. These omissions are non-essential details (e.g., greening ratio, floor area ratio) rather than key identifiers (e.g., buyer, area, quality). A second listing of the same parcel may thus exist simply to complete the initial record. The cases with extremely high numbers of duplicates appear to largely be erroneous; for instance, there are 73 duplicate entries of a parcel with an area of only 100 square meters purchased by a single buyer at the same price in the same area with the same usage, all marked as princelings. Another case with 71 apparent duplicates has a larger area, but in the online database, there is no record of even one transaction occurring with the given parcel features.
@@
Additional cases appear to be partially correct, partially incorrect duplicates. For instance, the Ministry of Land’s database reveals two identical parcels were sold at the same time to a single buyer at the same price. Yet, CK’s dataset has three transactions that match these features rather than two.777This may reflect an erroneous duplicate entry being removed from the website after CK obtained their data. A sample of several such cases is included in the Appendix.
@@
Further, some transactions appear identical except for the location dummy variables (near500 and near1500) created by CK. These are dummy variables that reflect whether a given parcel is within 500 or 1500 meters, respectively, of a parcel purchased by a princeling firm. Presumably, CK used the specific location data (not included in their replication files) to determine these binary values, and thus, having different binaries for these values implies that the parcels are indeed distinct. These rows are therefore not classified as duplicates in this paper.888There are approximately 100,000 of these rows in the data. I was unable to verify whether these parcels are distinct using the Ministry of Land’s database alone; the specific address information would be needed to ensure these transactions reflect separate parcels. Even if one wanted to remove these rows as duplicates, it is not apparent which observation contains the correct near500 and near1500 values without the parcel addresses. I thus leave these quasi-duplicates in, but note that their status as unique parcels should be verified further.
@@
Thus, a matched comparison with the Ministry of Land’s database, tracing parcel features and Transfer Announcements, suggests that a significant portion of the duplicate rows are erroneous, referring to the same underlying transaction. Yet, in the absence of specific location details for the parcels, it is not possible to examine the exact extent of incorrect duplicates versus those that refer to different but identical parcels.
@@ 2.1 Replicating results without duplicates: Tables V and X
When replicating CK’s regression results, I drop all duplicate rows but recognize that this path excludes duplicate entries that refer to different but identical parcels. However, as highlighted in Table 2, the regression results are largely still similar for their main specification (CK’s Table V, “Princeling Purchase and Land Price, 2004-2016”),
As detailed by CK, the dependent variable “is the log of price (yuan per square meter) for land parcel sold by local government to firm for usage in month-year . The key explanatory variable, , is a dummy variable equal to 1 if a firm is connected to a princeling” and 0 otherwise (p. 203). They control for several transaction-level variables including land quality (as evaluated by the municipal government, on a scale of 1-20), state-ownership status, sale method, firm size, and the log(area/100), where area is the land area sold in square meters. They also have city-year-usage, industry, and month fixed effects, with standard errors clustered at the province and firm levels and singletons included.
Column (1) illustrates regression (1) run on the full dataset, while columns (2) and (3) are conducted on the matched sample of parcels within a specified radius (1,500 or 500 meters, respectively) from a princeling parcel. Columns (4), (5), and (6) repeat these regressions but now include princeling firms interacted with the dummy variable (Politburo Standing Committee Member), aiming to explore whether firms connected to these committee members received larger discounts. Likewise, columns (7), (8), and (9) include princeling firms interacted with a dummy variable that reflects whether a transaction occurred after the connected official was retired. In all of these cases, the princeling effect is strongly negative with small standard errors—as was the case in CK’s original results—suggesting that the princeling effect is robust
@@
despite the apparent duplicates. Interestingly, when duplicates are removed, the princeling effect itself appears slightly smaller; for instance, the price advantage for princeling-connected firms, based on results for the 500-meter-radius sample, is now a 55.3% discount compared to CK’s original 57% discount. However, the interacted terms (princeling firms with Politburo Standing Committee connections and princeling firms connected to retired officials) are consistently larger than the original data’s results.999Even when replicating these main regression specifications on a subset without duplicates—and then going further to exclude rows which have different values in the near500 and near1500 binaries but are otherwise duplicates—the regression results shrink slightly more towards zero but remain largely the same. Standard errors likewise remain similar.
CK also use this granular transaction data in their Table X, “Princeling Purchases and Land Prices after Xi Took Office,” which examines whether President Xi’s anticorruption campaign impacted the price discount given to princeling firms. I regenerate this table without duplicate rows (Table 3), and the results are again quite similar to CK’s original findings: for the 500-meter matched sample, CK found that princeling firms had previously obtained an average price discount of 60.1% that dropped by 11.7% after the introduction of President Xi’s anticorruption
@@
campaign (column (2)); without duplicates, the 60.2% discount before the anticorruption campaign drops by 13.8% after the campaign begins (Table 3, column (2)). The other columns are equally similar and again suggest the anticorruption campaign strongly impacts the magnitude of discounts obtained by princelings; these remaining columns explore changes in princeling advantage in the presence of interacted dummy variables involving different dimensions of the anticorruption campaign (e.g., whether a central inspection occurred in the province at the time of sale, whether the transaction occurred after the anticorruption campaign began, etc.). As with CK’s Table V, these results are robust even in the presence of potentially erroneous duplicate rows.

### v2

# Are Princelings Truly Busted? Evaluating Transaction Discounts in China’s Land Market

@@ Abstract
This paper replicates Chen and Kung’s 2019 analysis (The Quarterly Journal of Economics 134(1): 185-226). Inspecting the data reveals that nearly one-third of transactions (388,903 out of 1,208,621) are perfect duplicates of other rows, excluding the transaction number. The analysis on the data sans duplicates replicates their statistically significant princeling effect, robust across various specifications. Further analysis reveals a disagreement between Chen and Kung’s text and code: the paper’s “logarithm of area” is actually area () divided by one million. This therefore necessitates a reinterpretation of the estimation results, revealing that the princeling effect is extremely large.
@@ 1 Introduction
The recent implosion of several leading Chinese real estate developers, such as Evergrande and Country Garden, has kindled interest in the mechanisms of the Chinese real estate market. Unique not only for its highly leveraged financing practices, the market is also a nexus of interactions between state-owned and private enterprises and has historically been plagued by corruption and cronyism, particularly on the land sale side. One often-cited paper on the topic is Chen and Kung (2019), titled “Busting the ‘Princelings’: The Campaign against Corruption in China’s Primary Land Market,” which explores the unique mechanism of Chinese cronyism in land sale. Specifically, they examine how “princeling” firms—those firms tied to family members of China’s top governing body, the Politburo—receive substantial price discounts on land parcels after controlling for several location-based and transaction-level variables ([2]). They also have significant findings relating princeling power, promotion likelihood, and magnitude of discount: more powerful princelings obtain larger discounts, and provincial party secretaries providing discounts were more likely to be promoted, with the likelihood increasing by the magnitude of the discount and the area of discounted land sold.
@@
Between containing detailed information about officials’ promotions and establishing firms’ state-ownership and princeling statuses, the paper’s replication data is also of popular interest and has over 12,200 downloads from its archive on the Harvard Dataverse.111The paper’s replication data can be found at https://doi.org/10.7910/DVN/XW6OJT.
@@
This paper replicates the results of Chen and Kung (2019, hereafter “CK”). A previous investigation by [5] diagnosed issues with CK’s variable on mayoral promotion. Yet, Wiebe found that the result was unchanged even when a corrected promotion variable was used. My replication identifies more extensive issues: an initial inspection of CK’s data reveals that nearly one-third of the transactions (388,903 out of 1,208,621) are perfect duplicates of other rows, excluding the transaction number. These duplicates suggest, for instance, that the same buyer purchased two (or more) identical parcels of land with the same quality rating, area, and usage in the same city, month, and year, at the same price. Examining records of land sale transactions suggests that some of these duplicates correctly refer to identical yet distinct parcels sold at the same time, but many other duplicates erroneously reference the same parcel of land multiple times. Importantly, over one-half of the princeling parcels in the full dataset (9,815 out of 19,812) are duplicates. Replicating the analysis on the data, sans duplicates, yields similar results to CK’s original findings: while the discount magnitude of princelings is slightly smaller than that found using the original dataset, the princeling effect is still large and statistically significant across the regression results.
@@
Yet, further inspection reveals that when CK aggregate parcel areas to group parcels (for instance, summing transaction data to calculate the total area of land sold to each firm in each year), their text and code differ: the “logarithm of area” in CK’s paper is actually area () divided by 1,000,000. I thus reinterpret CK’s originally calculated coefficients as scaled area values rather than logged area values, with and without duplicates. On this scale, the magnitude of the princeling effect is quite large, suggesting that princelings on average purchase tens of thousands of additional square meters of land per year compared to their nonprinceling peers—areas double, triple, or even quadruple the median parcel size, depending on the specification. In the appendix, I include calculations utilizing the logarithm of area purchased; these results are strange and somewhat unstable, with the princeling effect not only changing sign and direction between tables, but also dropping sharply prior to the anticorruption campaign.
@@ 2 Apparent Duplicates in the Data
In their paper, CK construct a set of “princeling” firms from several online sources and match them to land transaction data obtained from the Ministry of Land and Resources. I focus on this latter data, extracted from the Land Transaction Monitoring System (www.landchina.com). As each municipality’s “bureau of land and resources is required to report each land transaction in their jurisdiction electronically on this website,” per the Law of Land Management (p. 199), the dataset captures all land transactions.222China eliminated the secondary market for land use rights (LURs) in 2004, meaning the State is the sole seller of land, doing so at the municipal level ([4]; [3]). The Land Transaction Monitoring System provides comprehensive information on land transactions, relaying everything from transaction details like total payment amount, sale method, buyer name, and total area of land sold, to building restrictions, like floor area ratio and greening ratio, to precise location details on the parcel.333CK note that the system also provides the address and postal code of the parcel, “a two-digit code of land usage… [and] a three-digit industry code of the buyer’s firm” (p. 200). These fields are no longer included on the Land Transaction Monitoring System website nor are they included in CK’s replication data, which makes determining the extent of duplicates more complicated; it is possible that these information fields were removed from the database sometime after CK collected the data. Alternatively, CK may have obtained a slightly more comprehensive version of the website’s data directly from the Ministry.
@@
Examining CK’s dataset reveals several cases of apparent duplicates: specifically, after removing the transaction identifier, which was likely added by CK in data processing, 388,903 land transactions appear to be perfect duplicates, suggesting that a single buyer purchased identical parcels in the same month, year, and city for the same price and use. Table LABEL:tab:table_of_duplicates shows the extent and distribution of these duplicates. Of the 1,208,621 land transactions in CK’s data, 819,718 are unique, and 299,167 of these have at least one duplicate. While most transactions have only one duplicate, others have several—as many as 158. Importantly, of the 19,812 transactions that were conducted by princelings in the original data, over half of these are duplicates, with a significant share having three or more seemingly identical transactions. Duplicates are especially prevalent after 2012—the year that the anti-corruption campaign began.
| Total Observations | Observations by Time Period | ||||
| Frequency of occurrence | All | State-owned | Princeling | 2012 and earlier | After 2012 |
| 1 | 520,551 | 181,571 | 4,414 | 253,322 | 267,120 |
| 2 | 218,001 | 75,640 | 3,479 | 191,718 | 26,283 |
| 3 | 79,366 | 27,473 | 1,554 | 79,200 | 166 |
| 4 | 581 | 193 | 81 | 518 | 63 |
| 5-10 | 966 | 319 | 201 | 870 | 96 |
| 11-20 | 167 | 59 | 48 | 134 | 33 |
| 21-30 | 48 | 10 | 21 | 34 | 14 |
| 31-158 | 38 | 15 | 17 | 29 | 9 |
| Number of unique rows | 819,718 | 285,280 | 9,815 | 525,825 | 293,784 |
| Number of total rows in Chen and Kung | 1,208,621 | 420,004 | 19,812 | 886,116 | 322,396 |
| Number of duplicate rows | 388,903 | 134,724 | 9,997 | 360,291 | 28,612 |
| Number of unique rows with duplicates | 299,167 | 103,709 | 5,401 | 272,503 | 26,664 |
@@
In light of these results, there are a couple of possibilities: first, that there are truly identical neighboring parcels that are sold at the same time and price to a single buyer for specific usages. Second, it is possible that the duplicate rows are erroneous and that the database itself has multiple land transactions that refer to the same underlying land parcel. Examining CK’s replication data is not enough to confirm either possibility, as the data do not include each parcel’s specific address. Yet, the transaction details can be matched with the original data from the Ministry of Land’s database, and this additional information can be used to determine if multiple transaction listings refer to the same parcel. Specifically, after being identified in the database, each land transaction can be linked with a parcel number via the transaction’s original sale announcement (the “Transfer Announcement”), and one can confirm how many parcels of each size are up for sale in the area at the time. With this information, identical land transaction listings that refer to the same underlying parcel can be confirmed as erroneous duplicates.
@@
Looking at the detailed transaction records indicates that both hypotheses are true: while identical parcels were sometimes sold to a single buyer, many other duplicates refer to one underlying parcel and are thus erroneous. Some of these are quite easy to recognize as duplicate listings of the same land transaction when referencing the Ministry of Land’s database, as they may even have the same electronic lookup number—a “unique” identifier—for multiple transactions.444The electronic lookup number is the database’s main identifier/search key for unique land transactions. Having two “separate” transactions with the same electronic lookup number is thus impossible and indicates an error. The Transfer Announcements also confirm there should only be one listing for the parcel in the Ministry of Land’s database—and in CK’s data.
@@
Other cases are less clear but strongly indicative of incorrect duplicates. For instance, two transactions have different electronic lookup numbers, but the Transfer Announcement suggests that both listings refer to the same underlying parcel.555In many cases, examining the full transaction details of the two listings reveals missing information in one listing. These omissions are non-essential details (e.g., greening ratio, floor area ratio) rather than key identifiers (e.g., buyer, area, quality). A second listing of the same parcel may thus exist simply to complete the initial record. Cases with extremely high numbers of duplicates appear to largely be erroneous; for instance, for one entry with 71 apparent duplicates in CK’s data, there is no record of even one transaction occurring with the given parcel features.
@@
Additional cases appear to be partially correct, partially incorrect duplicates. For instance, the Ministry of Land’s database reveals two identical parcels were sold at the same time to a single buyer at the same price. Yet, CK’s dataset has three transactions that match these features rather than two.666This may reflect an erroneous duplicate entry being removed from the website after CK obtained their data. A sample of several such cases is included in the appendix.
@@
Further, some transactions appear identical except for the location dummy variables (near500 and near1500) created by CK. These are dummy variables that reflect whether a given parcel is within 500 or 1500 meters, respectively, of a parcel purchased by a princeling firm. Presumably, CK used the specific location data (not included in their replication files) to determine these binary values, and thus, having different binaries for these values implies that the parcels are indeed distinct. These rows are therefore not classified as duplicates in this paper.777There are approximately 100,000 of these rows in the data. I was unable to verify whether these parcels are distinct using the Ministry of Land’s database alone; the specific address information would be needed to ensure these transactions reflect separate parcels. Even if one wanted to remove these rows as duplicates, it is not apparent which observation contains the correct near500 and near1500 values without the parcel addresses. I thus leave these quasi-duplicates in, but note that their status as unique parcels should be verified further. I also conduct robustness checks that exclude these duplicates, as discussed in footnote 10.
@@
Thus, a matched comparison with the Ministry of Land’s database, tracing parcel features and Transfer Announcements, suggests a significant portion of the duplicate rows erroneously refer to the same underlying transaction. Yet, without parcel addresses, it is not possible to examine the exact extent of incorrect duplicates versus those that refer to different but identical parcels.
@@ 2.1 Replicating results without duplicates: Tables V and X
When replicating CK’s regression results, I drop all duplicate rows but recognize that this path excludes duplicate entries that refer to different but identical parcels. However, as highlighted in Table 2, the regression results are largely still similar for their main specification (CK’s Table V),
As detailed by CK, the dependent variable “is the log of price (yuan per square meter) for land parcel sold by local government to firm for usage in month-year . The key explanatory variable, , is a dummy variable equal to 1 if a firm is connected to a princeling” and 0 otherwise (p. 203). They control for several transaction-level variables including land quality (as evaluated by the municipal government, on a scale of 1-20), state-ownership status, sale method, firm size, and log(area/100), where area is the land area sold in square meters. They also have city-year-usage, industry, and month fixed effects, with standard errors clustered at the province and firm levels and singletons included.
Column (1) illustrates regression (1) run on the full dataset, while columns (2) and (3) are conducted on the matched sample of parcels within a specified radius (1,500 or 500 meters, respectively) from a princeling parcel. Columns (4), (5), and (6) repeat these regressions but now include princeling firms interacted with the dummy variable (Politburo Standing Committee Member), aiming to explore whether firms connected to these committee members received larger discounts. Likewise, columns (7), (8), and (9) include princeling firms interacted with a dummy variable that reflects whether a transaction occurred after the connected official was retired. In all of these cases, the princeling effect is strongly negative with small standard errors—as was the case in CK’s original results—suggesting that the princeling effect is robust despite the apparent duplicates. Interestingly, when duplicates are removed, the princeling effect itself appears slightly smaller; for instance, the price advantage for princeling-connected firms, based on results for the 500-meter-radius sample, is now a 55.3% (() discount compared to CK’s original 57.0% (() discount. However, the interacted terms (princeling firms with Politburo Standing Committee connections and princeling firms connected to retired officials) are consistently larger than the original data’s results.888As a robustness check, I also replicate these main regression specifications using a stronger definition of duplicates—that is, when defining a unique row, I exclude all variables added in data processing, like near500 and near1500. Under this definition, there are 727,127 unique rows. The regression results shrink slightly more towards zero but remain largely the same. Standard errors likewise remain similar.
@@
CK also use this granular transaction data in their Table X, “Princeling Purchases and Land Prices after Xi Took Office,” which examines whether President Xi’s anticorruption campaign impacted the price discount given to princeling firms. I regenerate this table without duplicate rows (Table 3), and the results are again quite similar to CK’s original findings: for the 500-meter matched sample, CK found that princeling firms had previously obtained an average price discount of 60.1% (() that dropped by 11.7% (() after the introduction of President Xi’s anticorruption campaign (column (2)); without duplicates, the 60.2% (() discount before the anticorruption campaign drops by 13.8% (() after the campaign begins (Table 3, column (2)). The other columns are equally similar and again suggest the anticorruption campaign strongly impacts the magnitude of discounts obtained by princelings; these remaining columns explore changes in princeling advantage in the presence of interacted dummy variables involving different dimensions of the anticorruption campaign (e.g.,
@@
whether a central inspection occurred in the province at the time of sale, whether the transaction occurred after the anticorruption campaign began, etc.). As with CK’s Table V, these results are robust even in the presence of potentially erroneous duplicate rows.